# Provably Efficient Online RLHF with One-Pass Reward Modeling

### v2

@@ 7.2.2 Active Data Collection
In this experimental setup, we constrain the algorithm to select only 6,400 samples from the full training dataset, based on different data selection strategies. We evaluate the performance of both the standard MLE-based method and our proposed OMD-based method under this limited data regime. To assess the effectiveness of the data selection strategy itself, we compare our approach against a random selection baseline. We report the results across four key metrics: (a) training loss curve, (b) evaluation accuracy curve and (c) final evaluation accuracy and training time. Figure 5 and Figure 6 demonstrate that our OMD-based method achieves competitive performance with the MLE-based method for both data collection strategies, while improving the training time by significantly. Moreover, our data selection strategy outperforms the random selection strategy, demonstrating that our method can effectively select informative data to improve the performance.
7.2.3 Deployment-Time Adaptation
To emulate real-world deployment scenarios, we partition the dataset into 20 sequential chunks and process them incrementally. This setup allows us to evaluate the adaptability and robustness of different action selection strategies over time. We benchmark our proposed selection strategy against three baselines: (i): random selection, (ii): the best and second best actions, and (iii): the best and worst actions. We combine the above strategies with MLE-based and OMD-based reward estimators. We report three key metrics: (a) the average cumulative reward for MLE-based methods, (b) the average cumulative reward for OMD-based methods, and (c) the win rate of OMD-based methods over MLE-based counterparts under identical selection strategies. As illustrated in Figure 7 and Figure 8, our action selection strategy consistently outperforms the baselines across both estimator types. These results highlight the strength of our approach in balancing the exploitation of high-reward responses with sufficient exploration to drive continual model improvement. Moreover, the win rate comparisons suggest that our OMD-based estimator shows competitive performance relative to its MLE-based counterpart, further validating its practical effectiveness.
(a) Reward of MLE-based methods
(b) Reward of OMD-based methods
@@ 7.2.3 Deployment-Time Adaptation
Figure 7: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
(a) Reward of MLE-based methods
(b) Reward of OMD-based methods
(c) Win rates, all methods
Figure 8: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
@@ 8 Conclusion
In this work, we study the problem of online RLHF from the perspective of contextual bandits. To address the linear growth in computational complexity with respect to the number of iterations in reward modeling, we propose a novel one-pass reward modeling algorithm that achieves constant-time complexity per iteration. Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information. We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation, and design tailored algorithms for each scenario. We provide theoretical guarantees that our approach achieves superior statistical and computational efficiency relative to existing methods. Finally, we provide practical implementations of our method and conduct experiments using the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, empirically validating its practical effectiveness.
@@ Appendix A Useful Lemmas
For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$ Then, for the following update rule
@@
[…]
@@
~$\big{\langle}\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\right)$
@@
$\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}=$ $\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\ell_{i}(\widetilde{\theta}_{i})-\nabla^{2}\ell_{i}(\widetilde{\theta}_{i})(\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
=\ = $\big{\langle}D^{3}\ell_{i}({\xi}_{i+1})[\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}](\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
@@
[…]
@@ Appendix B Proof of Lemma 1
[…]
It remains to bound the right-hand side of the above inequality in the following. The most challenging part is to bound the term $\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})$. This term might seem straightforward to control, as it can be observed that $\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}(\theta)\triangleq\mathbb{E}_{y_{i}}[\ell_{i}(\theta)]$, where $\ell_{i}(\theta)$ serves as an empirical observation of ¯$\bar{\ell}(\theta)$. Consequently, the loss gap term seemingly can be bounded using appropriate concentration results. However, a caveat lies in the fact that the update of the estimator $\widetilde{\theta}_{i+1}$ depends on $\ell_{i}$, or more precisely $y_{i}$, making it difficult to directly apply such concentrations.
To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as $[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}$ and $[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)}$, where $[\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
@@
Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with $[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))$ and $[\sigma^{-1}(p)]_{0}=\log((1-p)/{p})$. Then, we decompose the regret into two terms by introducing an intermediate term.
@@
$\mathtt{term~{}(a)}\leq\left(3\log(1+2t)+2+LB\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2.$
@@
where $C=2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$. Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
@@
[…]
@@
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}\leq\mathcal{O}\Big{(}\sqrt{d}(\log(t/\delta))^{2}\Big{)}.$
@@ Appendix C Proof of Theorem 1
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)=$ $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}\phi(s,\pi^{*}(s))\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi^{*}(s))\right]$
=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\phi(x,\pi^{*}(x))\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
@@
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (13)
@@
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
@@ Appendix E Proof of Theorem 3
=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$
@@
[…]
@@
[…]
@@
[…]
@@
=\ = $3\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
where the first inequality holds by the Holder’s inequality and the arm selection strategy of $a_{t}$ such that $\phi(x_{t},\pi^{*}(x_{t}))^{\top}\widetilde{\theta}_{t}\leq\phi(x_{t},a_{t})^{\top}\widetilde{\theta}_{t}$, the second inequality holds by $\widetilde{\theta}_{t}\in\mathcal{C}_{t}$ with probability at least $1-\delta$ by Lemma 1, the third inequality holds by arm selection strategy of $a_{t}^{\prime}$ such that $a_{t}^{\prime}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widetilde{\theta}_{t}+2\widetilde{\beta}\lVert\phi(x_{t},a)-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$.

### v3

@@ 7.2 Experimental results
Active data collection.  In this setting, we only allow the algorithm to select $6,400$ samples out of the whole training datasets for training according to different selection strategies. To evaluate the effectiveness of the data selection strategy, we compare our method with the random selection strategy. We evaluate the performance of the MLE-based method and our proposed OMD-based method. Figure 3 demonstrates that our OMD-based method achieves comparable performance with the MLE-based method for both data collection strategies, while improving the training time by approximately three times. Moreover, our data selection strategy outperforms the random selection strategy, showing that our method can effectively select informative data to improve the performance.
Deployment-time adaptation.  We divide the dataset into 20 chunks and process them sequentially to simulate the deployment scenario. We compare our action selection strategy with (i): random selection, (ii): select the best and second best actions, and (iii): select the best and worst actions. We combine the above strategies with MLE-based and OMD-based methods. We report both the average cumulative rewards and win rates of each method, where the win rate is defined as the proportion of pairwise comparisons in which a method outperforms all others. As shown in Figure 4, our action selection strategy outperforms the baselines for both MLE-based and OMD-based methods. This validates the effectiveness of our selection strategy that balances the exploitation of high-reward responses with sufficient exploration to facilitate model improvement. Besides, the win rates show that our OMD-based method achieves competitive performance with the MLE-based method.
(a) Rewards of MLE methods
(b) Rewards of OMD methods
@@
Figure 4: For online RLHF with deployment-time adaptation, we report (a) cumulative rewards of MLE-based methods, (b) cumulative rewards of OMD-based methods, and (c) win rates.
@@ 8 Conclusion
In this work, we address a key challenge in online RLHF, where the computational complexity typically grows linearly with the number of iterations. To overcome this limitation, we propose a novel one-pass algorithm that eliminates the need to store historical data and achieves constant-time complexity per iteration. Our approach is built upon the online mirror descent framework with a carefully designed local norm. We apply our method to three online RLHF settings and design tailored algorithms for each scenario. We provide both theoretical guarantees and efficient implementations, demonstrating that our approach improves statistical and computational efficiency over existing methods. Finally, we validate the effectiveness of our method through extensive experiments.
@@
Acknowledgments
This research was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382, 62206125). Peng Zhao was supported in part by the Xiaomi Foundation.
@@ Appendix A Useful Lemmas
For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as
$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$
Then, for the following update rule
@@
$\leq{}$ $2\eta\left(\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})\right)+4\lambda B^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{2}^{2}-\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{\mathcal{H}_{i}}^{2}.$
@@
$\big\langle\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\right)$
@@
$\big\langle\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle=\$$\big\langle\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\ell_{i}(\widetilde{\theta}_{i})-\nabla^{2}\ell_{i}(\widetilde{\theta}_{i})(\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle$
@@
$\leqslant\$$3\sqrt{2}L\big\|\widetilde{\theta}_{i+1}-\theta^{*}\big\|_{2}\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{\nabla^{2}\ell_{i}({\xi}_{i+1})}^{2}$
$\leqslant\$$6\sqrt{2}BL\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{\nabla^{2}\ell_{i}({\xi}_{i+1})}^{2}$
$\leqslant\$$6\sqrt{2}BL^{3}\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{2}^{2},$
@@ Appendix B Proof of Lemma 1
$\leq{}$ $2\eta\left(\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})\right)+4\lambda B^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{2}^{2}-\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{\mathcal{H}_{i}}^{2}.$
It remains to bound the right-hand side of the above inequality in the following. The most challenging part is to bound the term $\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})$. This term might seem straightforward to control, as it can be observed that $\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}(\theta)\triangleq\mathbb{E}_{y_{i}}[\ell_{i}(\theta)]$, where $\ell_{i}(\theta)$ serves as an empirical observation of $\bar{\ell}(\theta)$. Consequently, the loss gap term seemingly can be bounded using appropriate concentration results. However, a caveat lies in the fact that the update of the estimator $\widetilde{\theta}_{i+1}$ depends on $\ell_{i}$, or more precisely $y_{i}$, making it difficult to directly apply such concentrations.
To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as
$[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}\quad\text{and}\quad[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)},$
where $\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
@@
Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with
$[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))\quad\text{and}\quad[\sigma^{-1}(p)]_{0}=\log((1-p)/{p}).$
Then, we decompose the regret into two terms by introducing an intermediate term.
@@
$\mathtt{term~(a)}\leq 11\cdot(3\log(1+2t)+2+LB)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2.$
@@
where $C=22\eta(3\log(1+2t)+2+LB)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$. Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
@@
$\leq\$$\left(12\sqrt{2}BL^{3}\eta-\frac{\lambda}{7}\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}$
$\leq\$$0.$
@@
$\big\|\widetilde{\theta}_{t+1}-\theta^{*}\big\|_{\mathcal{H}_{t+1}}\leq\mathcal{O}\Big(\sqrt{d}\log(t/\delta)\Big).$
@@ Appendix C Proof of Theorem 1
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)$
@@
[…]
$\leq\$$\Big(\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}+\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\Big)\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}},$
@@
$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq\$$2\widetilde{\beta}_{T}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}.$ (10)
@@
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)$ $\leq 2\widetilde{\beta}_{T}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}$
$\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}\right).$
@@ Appendix E Proof of Theorem 3
$=\$$\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\theta^{*}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\theta^{*}$
$=\$$2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\theta^{*}+\big(\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\theta^{*}$
$=\$$2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\widetilde{\theta}_{t}$
@@
$\leq\$$2\big\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}\big\lVert\theta^{*}-\widetilde{\theta}_{t}\big\rVert_{\mathcal{H}_{t}}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\widetilde{\theta}_{t}$
@@
$\leq\$$2\widetilde{\beta}_{t}\big\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\widetilde{\theta}_{t}$
@@
$\leq\$$2\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t}^{\prime})-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}+(\phi(x_{t},a_{t}^{\prime})-\phi(x_{t},\pi^{*}(x_{t})))^{\top}\widetilde{\theta}_{t}$
@@
$=\$$3\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}},$
where the first inequality holds by the Holder’s inequality and the arm selection strategy of $a_{t}$ such that $\phi(x_{t},\pi^{*}(x_{t}))^{\top}\widetilde{\theta}_{t}\leq\phi(x_{t},a_{t})^{\top}\widetilde{\theta}_{t}$, the second inequality holds by $\widetilde{\theta}_{t}\in\mathcal{C}_{t}$ with probability at least $1-\delta$ by Lemma 1, the third inequality holds by arm selection strategy of $a_{t}^{\prime}$ such that $a_{t}^{\prime}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widetilde{\theta}_{t}+2\widetilde{\beta}\lVert\phi(x_{t},a)-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$. Thus, we have
${\operatorname*{Reg}}_{T}$ $=\frac{1}{2}\sum_{t=1}^{T}\Big(\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big)+\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big)\Big)$
$\leq\frac{3}{2}\widetilde{\beta}_{T}\sum_{t=1}^{T}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}}.$